1. A population has a mean of 129. If a random sample of 8 items from the population results in the following sampled​ values, what is the sampling error for the​ sample? 145, 129, 108, 153, 149, 154, 117, 120

2. For a population with a mean equal to 250 and a standard deviation equal to 25​, calculate the standard error of the mean for the following sample sizes.

​a) 10 ​b) 40 ​c) 70

3. Managers at a local phone service wireless retail center have a goal that 74​% of the​ center's customers will have to wait less than five minutes for service. The data below shows the wait times of a random sample of 25 customers in minutes. Does this sample provide evidence that​ management's goal is being​ achieved?

4.3 5.6 2.4 10.4 3.7
3.6 0.3 7.9 4.7 4.5
4.9 2.3 3.2 9.3 6.8
5.4 4.3 0.3 2.8 5.6
5.9 3.2 2.3 2.7 4.2

4. A college has 250 full-time employees that are currently covered under the​ school's health care plan. The average​ out-of-pocket cost for the employees on the plan is $1,820 with a standard deviation of $515. The college is performing an audit of its health care plan and has randomly selected 30 employees to analyze their​ out-of-pocket costs.

a. Calculate the standard error of the mean.

b. What is the probability that the sample mean will be less than $1,765​?

c. What is the probability that the sample mean will be more than $1,800​?

d. What is the probability that the sample mean will be between $1,845 and $1,890​?

5. For a population with a proportion equal to 0.28​, calculate the standard error of the proportion for the following sample sizes. a) 50 b) 100 c)150
Answer #1

1) Mean of the sample = (145+129+108+153+149+154+117+120) / 8 = 1075/8 = 134.375

Sampling error = sample mean - population mean = 134.375 - 129 =5.375